if 4 sweaters cost $51.04, how much would 6 sweaters cost?

Found 2 solutions by rfer, tutorcecilia:
Answer by rfer(16322)

4/51.04=6/x
4x=306.24
x=$76.56

Answer by tutorcecilia(2152)

This is a problem involving ratios. First find the ratio of dollars to sweaters. That means, how much would just 1 sweater cost if 4 sweaters cost $51.04?
.
if 4 sweaters cost $51.04:
$51.04 / 4 sweaters = $12.76 / 1 sweater
.
Since one sweater costs $12.76, multiply this by 6 sweaters and you get:
.
how much would 6 sweaters cost?
$12.76 x 6 sweaters = $76.56